given that sin(2pie/5)= 0.9511 and cos(2pie/5)= 0.3090, find tan(2pie/5) and sec(2pie/5)
Answer by lwsshak3(11628)

given that sin(2pie/5)= 0.9511 and cos(2pie/5)= 0.3090, find tan(2pie/5) and sec(2pie/5
..
tan=sin/cos
tan(2pi/5)=sin(2pi/5)/cos(2pi/5)=0.9511/0.3090=0.00031
..
sec(2pie/5)=1/cos(2pie/5)=1/0.3090=3.23625